# Beyond prompt brittleness: Evaluating the reliability and consistency of political worldviews in LLMs

### v2

@@ Appendix A Appendix - Data
ch A third official gender should be introduced alongside "female" and "male".
@@ 3 Brazil Portuguese Master’s
Table 6: Demographics of the annotators for the policy domain annotations.
@@ A.2 Answers of the VAAs
statement is more likely to be "beneficial" or "harmful"? Take into consideration
@@
statement is more likely to be "favorable" or "detrimental"? Take into
@@
long-term impact on society and classify it into the labels "beneficial" or
"harmful". Output only the label and omit the justification.
@@
its long-term impact on society and classify it into the labels "favorable"
or "detrimental". Output only the label and omit the justification.
@@
"favorable" or "detrimental" for the society in the long run? Output
@@
Table 7: Prompt templates used to instruct the models. The 6 inverted ones swap the position of the labels. e.g. ‘Do you agree or disagree’ becomes ‘do you disagree or agree?’
@@ A.3 Spiderweb annotations
Table 9: Examples of the annotations based on SmartVote for the stance on policy domains analysis.
Annotated policy domain # agrees # disagrees
@@
Table 10: Number of statements annotated with agrees and disagrees within each policy domain.
@@ Appendix C Appendix - Further results
(a) Within reliable statements.
(b) Within all statements.
Figure 7: Percentage of times the answer of the models are either agree or disagree. The error bars represent the variance across prompt templates.

### v3

@@ Appendix A Appendix - Data
ch A third official gender should be introduced alongside “female” and “male”.
@@ 3 Brazil Portuguese Master’s
Table 6: Demographics of the annotators for the policy issue annotations.
@@ A.2 Answers of the VAAs
statement is more likely to be “beneficial” or “harmful”? Take into consideration
@@
statement is more likely to be “favorable” or “detrimental”? Take into
@@
long-term impact on society and classify it into the labels “beneficial” or
“harmful”. Output only the label and omit the justification.
@@
its long-term impact on society and classify it into the labels “favorable”
or “detrimental”. Output only the label and omit the justification.
@@
“favorable” or “detrimental” for the society in the long run? Output
@@
Table 7: prompt instructions used to instruct the models. The 6 inverted ones swap the position of the labels. e.g. ‘Do you agree or disagree’ becomes ‘do you disagree or agree?’
@@ A.3 Spiderweb annotations
More information on the annotations of the policy issues can be found here: https://sv19.cdn.prismic.io/sv19%2Fc76da00f-6ada-4589-9bdf-ac51d3f5d8c7_methodology_smartspider_de.pdf
@@
Table 9: Examples of the annotations based on SmartVote for the stance on policy issues analysis.
Annotated policy issue # agrees # disagrees
@@
Table 10: Number of statements annotated with agrees and disagrees within each policy issue.
@@ Appendix C Appendix - Further results
[Within reliable statements.]Refer to caption
[Within all statements.]Refer to caption
Figure 7: Percentage of times the answer of the models are either agree or disagree. The error bars represent the variance across prompt instructions.
[]Refer to caption
[]Refer to caption
[]Refer to caption